### v2

# RichRAG: Crafting Rich Responses for Multi-faceted Queries in Retrieval-Augmented Generation

Abstract
Retrieval-augmented generation (RAG) effectively addresses issues of static knowledge and hallucination in large language models. Existing studies mostly focus on question scenarios with clear user intents and concise answers. However, it is prevalent that users issue broad, open-ended queries with diverse sub-intents, for which they desire rich and long-form answers covering multiple relevant aspects. To tackle this important yet underexplored problem, we propose a novel RAG framework, namely RichRAG. It includes a sub-aspect explorer to identify potential sub-aspects of input questions, a multi-faceted retriever to build a candidate pool of diverse external documents related to these sub-aspects, and a generative list-wise ranker, which is a key module to provide the top-k most valuable documents for the final generator. These ranked documents sufficiently cover various query aspects and are aware of the generator’s preferences, hence incentivizing it to produce rich and comprehensive responses for users. The training of our ranker involves a supervised fine-tuning stage to ensure the basic coverage of documents, and a reinforcement learning stage to align downstream LLM’s preferences to the ranking of documents. Experimental results on two publicly available datasets prove that our framework effectively and efficiently provides comprehensive and satisfying responses to users.
RichRAG: Crafting Rich Responses for Multi-faceted Queries in Retrieval-Augmented Generation
Shuting Wang1, Xin Yu2, Mang Wang2, Weipeng Chen2, Yutao Zhu1and Zhicheng Dou1∗ 1Gaoling School of Artificial Intelligence, Renmin University of China 2Baichuan Intelligent Technology {wangshuting, dou}@ruc.edu.cn
**footnotetext: Corresponding author.
1 Introduction
Figure 1: An example of a scenario where a multi-faceted query requires a comprehensive answer.
Large language models (LLMs) have revolutionized how information is accessed online, shifting from returning ranked lists of relevant documents to directly generating answers to user queries. However, they still suffer from hallucinations and information staleness issues, impacting the authenticity and reliability of generated answers. Retrieval-augmented generation (RAG) has emerged as a promising solution, empowering LLMs to leverage reliable information from retrieved documents, thereby returning more reliable responses.
Though some advanced techniques Jiang et al. (2023); Asai et al. (2024); Wang et al. (2023c); Li et al. (2024) have been proposed, existing studies primarily focus on addressing specific problems that require concise and definitive answers. However, user intents are complex and multi-faceted, necessitating rich and comprehensive answers. As Figure 1 shows, when a user inquires about rapping-related information, a rich response about various aspects of rapping, such as origins, characteristics, and evolution could lead to a more satisfactory user experience than a superficial description.
Our research is focused on developing effective RAG approaches to handle these more complex user needs. We propose a RAG framework, RichRAG, which is designed to offer diverse external knowledge that comprehensively covers various sub-aspects of multi-faceted queries, thereby enhancing the downstream generator (an LLM) to yield rich responses. RichRAG first employs a sub-aspect explorer to explicitly predict sub-aspects of queries. Then, it adopts a multi-facet retriever to build a broad pool of candidate documents covering those identified sub-aspects. However, such redundant candidates inevitably contain much irrelevant noise and are hard to handle completely by LLMs due to limited input length. As a result, sorting out the top-k most valuable documents from the candidate pool is critical to the success of the RichRAG framework.
In further, we claim that a promising top-k ranking should have the following desirable features: (1) Comprehensiveness. Incentivizing LLM to generate rich and reliable responses requires external documents to comprehensively cover various query aspects. Therefore, the ability to model relationships among documents, hence maximizing the coverage of the entire reference list is crucial for the ranking module. (2) Alignment with the LLMs’ preferences. In RAG systems, the users of IR models are LLMs instead of humans. Thus, the reference order should be LLM-friendly, hence enhancing the generator to produce satisfying responses.
To achieve this, we devise a generative list-wise ranker based on encoder-decoder structures. It takes as input the user query, its identified sub-aspects, and all candidates, then directly generates top-k document IDs as final ranking lists. This structure offers two key advantages: (1) Global Document Modeling. The seq-to-seq model structure equips the ranker to effectively model global interactions among candidate documents, queries, and sub-aspects, thereby capturing the overall utility of generated ranking lists in covering the query’s multi-aspects. (2) Efficiency. Following the FiD structure Izacard and Grave (2021), we parallelize the encoding of each candidate and further introduce pooling and reuse operations to the decoder module. These strategies significantly reduce the spatiotemporal load of the ranker.
The optimization of our ranker consists of two stages: The first is supervised fine-tuning (SFT). To enhance the coverage of generated ranking lists on query aspects, we devise a coverage utility function based on which to build silver generation targets for training samples greedily. These silver targets allow us to SFT our ranker and ensure its basic ability. To further improve the ranking quality and align ranking lists with LLMs’ preferences, a reinforcement learning stage is introduced. We measure the LLMs’ final response quality to create reward values and adopt the DPO Rafailov et al. (2023) algorithm to optimize our ranker. In addition, we devise a unilateral significance sampling strategy (US3) to build valuable training samples for stable optimization. Experiments on two public datasets prove that RichRAG can effectively and efficiently generate more comprehensive answers for multi-faceted queries than existing methods.
Our contributions are three-fold:
(1) We propose a RAG framework RichRAG to explicitly model the query’s various sub-aspects, thereby providing comprehensive long-form responses to satisfy the user’s rich intents.
(2) We develop an efficient generative list-wise ranker that models the global gain of ranking lists considering rich user intents, delivering promising ranking lists for downstream LLMs.
(3) We devise the US3 approach to create reliable and valuable training pairs for the DPO algorithm, improving the quality and stability of optimization.
2 Related Works
2.1 Retrieval-augmented Generation
The RAG paradigm has exhibited excellent abilities in enhancing LLMs to generate accurate responses. To effectively optimize the system, some studies focus on concurrently or asynchronously training retrievers and generators Izacard et al. (2024); Borgeaud et al. (2022); Wang et al. (2023a); Arora et al. (2023); Paranjape et al. (2022); Lewis et al. (2020); Lin et al. (2023); Asai et al. (2024). However, as LLMs grow in size and inference costs rise, researchers have also explored fixing LLMs and optimizing retrievers as plug-in modules Shi et al. (2023); Yu et al. (2023) or introducing post-retrieval components, e.g., compressors and rankers Xu et al. (2023); Wang et al. (2023c); Li et al. (2024); Ke et al. (2024); Gao et al. (2024). Some studies Chan et al. (2024); Wang et al. (2024a); Khot et al. (2023); Yao et al. (2023) propose to decompose multi-hop questions into sub-tasks and solve them step-by-step to produce specific answers. However, our study mainly focuses on the query scenarios with various implicit sub-intents that require rich and comprehensive responses. It is different from the query decomposition scenario. Recent studies Rackauckas (2024) also highlight the importance of exploring users’ sub-intents, but its simplistic pipeline fails to model global document-intent interactions, leading to sub-optimal results.
2.2 Generative Ranking with LLM
Recently, the rise of LLMs allows researchers to establish various generative ranking models Nogueira et al. (2020); Zhuang et al. (2023a); Sun et al. (2023); Tamber et al. (2023); Zhuang et al. (2023b); Qin et al. (2024), including point-wise Nogueira et al. (2020); Zhuang et al. (2023a), pair-wise Zhuang et al. (2023b); Qin et al. (2024), and list-wise Tamber et al. (2023); Sun et al. (2023) models. To handle extensive document load, some generative list-wise methods Sun et al. (2023); Tamber et al. (2023) adopt a sliding-window approach to iteratively generate final ranking lists. Some studies Ke et al. (2024); Xu et al. (2024) also explore list-wise rankers in RAG systems but still focus on scenarios with specific intents and answers, neglecting the depth and breadth of user questions.
3 Method
In this section, we demonstrate our RichRAG framework, which explicitly considers the sub-aspects of multi-faceted questions to provide diverse and LLM-friendly external reference lists, thereby enhancing the richness and satisfaction of generated responses. Figure 2 (a) displays the overall framework. We first define the problem, and then delve into the introduction of each component, including the sub-aspect explorer, the multi-faceted retriever, and the generative list-wise ranker.
Figure 2: The overall framework of RichRAG. We describe the training stages of our ranker at the bottom.
3.1 Problem Definition
The basic RAG setting usually contains a knowledge corpus $\mathcal{C}$, a fixed retriever $\mathcal{R}$, and a fixed LLM serving as the generator, $\mathcal{G}$. For a multi-faceted query, $q$, its various subordinate aspects are denoted as $\mathcal{S}=\{s_{1},\ldots,s_{n}\}$. These sub-aspects have corresponding sub-answers, which are denoted as $\mathcal{A}=\{a_{1},\ldots,a_{n}\}$. [^1] The collection of sub-aspects and sub-answers is introduced in Appendix A. The combination of these sub-answers forms the ground truth answer, $a$, which is long-form and responds to all sub-aspects. Existing RAG models primarily focus on retrieving relevant documents from the corpus and incorporating them into the LLM’s input to generate responses closely aligned with ground truth answers. In this study, we aim to make responses, $r$, generated by RichRAG not only match the ground truth answers but also sufficiently cover individual sub-answers comprehensively, to ensure the responses’ richness and completeness.
3.2 Sub-aspect Explorer
Examining various sub-aspects under a user’s query could provide explicit insights into the user’s underlying intents, thereby enabling more satisfactory results for users Santos et al. (2010); Dang and Croft (2012); Liu et al. (2020); Wang et al. (2023b). We leverage LLMs to build our sub-aspect explorer, $\mathcal{E}$, due to their extensive world knowledge and excellent capabilities in language understanding and generation. This module takes a prompt $p_{se}$, which instructs the LLM to predict the sub-aspects of the input query, and a user’s query, $q$, as input and generates a series of sub-aspects under the query, i.e., ${\hat{\mathcal{S}}}=\{\hat{s}_{1},\ldots,\hat{s}_{m}\}$, as follows:
$\{\hat{s}_{1},\ldots,\hat{s}_{m}\}=\mathcal{E}(q,p_{se}).$ (1)
To align the sub-aspect explorer with the output format and the distribution of downstream data, we fine-tune it using training queries and their labeled sub-aspects. The target output is a concatenation of labeled sub-aspects surrounded by square brackets: $o=[s_{1}]\ldots[s_{n}].$ Subsequently, we optimize the sub-aspect explorer by the next token prediction (NTP) loss function:
$\mathcal{L}_{se}=-\sum\nolimits_{i=1}^{|o|}\log P(o_{i}|o_{1:i-1},q,p_{se}).$ (2)
3.3 Multi-faceted Retriever
Given the query’s sub-aspects that represent the user’s various potential sub-intents, we then use a multi-faceted retriever to collect documents that are relevant to various sub-aspects to build a diverse candidate pool. This operation could filter out apparent irrelevant documents and shrink the search space of the subsequent ranker. The multi-faceted retriever consists of the following two processes.
The first is a retrieval process, where we separately retrieve top-N documents $\mathcal{D}_{i}$ from the corpus for each sub-aspect $\hat{s_{i}}$. To avoid the topic drift, we concatenate each sub-aspect with the original query to form a new query and retrieve as follows,
$\mathcal{D}_{i}=\mathcal{R}(q\circ\hat{s_{i}},\mathcal{C}),$ (3)
where $\circ$ denotes concatenation and each document in $\mathcal{D}_{i}$ is associated with the sub-aspect, $\hat{s_{i}}$.
Next, a combination process is introduced to merge all these retrieved documents to create the candidate pool, $\mathcal{P}$. Since some documents may be retrieved multiple times by different sub-aspects, to reduce the space-time burden of the ranker, we treat the repeated documents as a single one, hence the associated sub-aspects form a set, $s(d)$:
$\mathcal{P}$ $=\mathcal{M}(\mathcal{D}_{1:m})=\{d_{1},\ldots,d_{M}\},$ (4)
$s(d_{i})$ =$=\{\hat{s}^{i}_{1},\ldots,\hat{s}^{i}_{n_{i}}\}.$
$\mathcal{M}()$ denotes the combination and $M$ is the maximum capacity of the pool. This candidate pool collects potentially valuable documents that sufficiently cover various sub-aspects of the query.
3.4 Generative List-wise Ranker
Though we have collected plenty of candidates related to various sub-aspects, directly providing these massive documents to the generator is challenging due to the extensive processing burden and potential noisy information. Consequently, we devise a ranking model that targets to sort out the top-$k$ most valuable documents from the candidate pool. These ranked documents should collectively cover the query’s various sub-aspects and adhere to the preferences of the generator, hence enhancing the response performance. To equip our ranker with the ability to globally model relationships among candidates, we build it upon a generative model, T5 Raffel et al. (2020), which views all candidates, sub-aspects, and the query as input and directly generates a top-k ranking list of document IDs (docids). For each candidate, $d_{i}$, we concatenate it with the original query, $q$, its associated sub-aspects, $s(d_{i})=\{\hat{s}^{i}_{1},\ldots,\hat{s}^{i}_{n_{i}}\}$, and some special tokens to formulate an input sequence:
$I_{i}=[\mathrm{D}_{i}]\circ q\circ[\mathrm{Q}]\circ\hat{s}^{i}_{1}\circ[\mathrm{E}]\ldots\;\hat{s}^{i}_{n_{i}}\circ[\mathrm{S}]\circ d_{i},$ (5)
where $[\mathrm{D}_{i}]$ is the docid token indicating the $i$-th candidate, $[\mathrm{Q}]$ denotes the end of the query, $[\mathrm{E}]$ separates each sub-aspect, and $[\mathrm{S}]$ denotes the end of associated sub-aspects. Inspired by FiD structure Izacard and Grave (2021), the encoder module, $\mathrm{Enc}()$, encodes candidate sequences in parallel to ensure high efficiency. Furthermore, since the ranker’s generation space is limited to docid tokens instead of the whole vocabulary, we use the pooling operation $\mathrm{Pool}()$ to extract the encoded output of docid tokens $\mathrm{\mathbf{e}}_{i}$ as relevance representations of candidates. They are then connected and entered into the decoder $\mathrm{Dec}()$ to generate the ranking list:
$,\ldots,[\mathrm{D}_{r(k)}]$ $=\mathrm{Dec}([\mathrm{\mathbf{e}}_{1};\ldots;\mathrm{\mathbf{e}}_{M}]),$ (6)
$\mathrm{\mathbf{e}}_{i}$ $=\mathrm{Pool}(\mathrm{Enc}(I_{i})),$
where $r(i)$ project the rank position $i$ into the generated docid that is ranked at the $i$-th position. This pooling operation could significantly reduce the time-space burden of the decoder.
Additionally, we implement a reuse strategy on the language model (LM) head layer to reduce unnecessary load and enhance the modeling accuracy. It sets the LM head layer’s projection matrix $\mathrm{\mathbf{F}}$, which maps generated hidden states to probabilistic token spaces, to be $d\times M$, ($M$ is the maximum number of candidates). Furthermore, our preliminary experiments imply that randomly initializing the value of $\mathrm{\mathbf{F}}$ is hard to optimize due to limited training samples. Therefore, we define its value using relevance representations of candidates to simplify optimization difficulty, hence improving the ranking performance. Thus, the probability of the $t$-th token is computed as below,
$p^{t}(d)=\mathrm{Softmax}\left(\mathbbm{M}(\mathrm{\mathbf{h}}^{t}\mathrm{\mathbf{F}}/{\tau})\right),\mathrm{\mathbf{F}}=[\mathrm{\mathbf{e}}_{1};\ldots;\mathrm{\mathbf{e}}_{M}],$ (7)
where $\mathrm{\mathbf{h}}^{t}$ is generated hidden states of the $t$-th token and $\tau$ is temperature to control the sharpness of distribution. $\mathbbm{M}()$ denotes a masking mechanism to set the probabilities of previously generated documents to zero, avoiding repetition. To ensure the ranker’s performance, we employ a two-stage optimization. We demonstrate it in the following sections and visualize it in Figure 2 (b).
3.4.1 Supervised Fine-tuning
To address the problem of the vast searching space of possible permutation, a major challenge for list-wise ranking algorithms, we adopt a greedy algorithm to build silver target ranking lists for each training instance, supporting supervised fine-tuning of the ranker. Specifically, we devise a coverage utility function, $\Phi^{t}(d)$, to measure the incremental gain in aspect coverage for each remaining document, $d$, conditioned on previously selected ones, $L^{*}_{t-1}$. The greedy selection is presented as follows,
$L^{*}_{t}$ $=L^{*}_{t-1}\cup d^{*}_{t},\;\;d^{*}_{t}=\mathop{\arg\max}\nolimits_{d\in\mathcal{P}/L^{*}_{t-1}}\Phi^{t}(d),$ (8)
$\Phi^{t}(d)$ $=\sum\nolimits_{i=1}^{n}w^{t}_{i}\phi(d,a_{i}).$
$\Phi^{t}(d)$ considers the current importance of each sub-aspect $w^{t}_{i}$ and the candidate’s coverage for each sub-aspect $\phi(d,a_{i})$. The coverage function $\phi()$ is implemented by the rouge-score between $d$ and the $i$-th sub-answer $a_{i}$. The current importance of sub-aspects is measured by calculating their coverage by previous $t-1$ selected documents using the following function with sum normalization $\mathrm{Norm}()$:
$w^{t}_{i}=1-\mathrm{Norm}_{i}({\max\nolimits_{\tilde{d}\in L^{*}_{t-1}}\phi(\tilde{d},a_{i})}),$ (9)
The silver target list, $L^{*}_{k}$, allows us to supervised fine-tune the ranker via the NTP task:
$\mathcal{L}_{sft}=-\sum\nolimits_{t=1}^{k}\log p(d^{*}_{t}|q,\mathcal{P},\hat{\mathcal{S}},L^{*}_{t-1}).$ (10)
3.4.2 Reinforcement Learning
After supervised fine-tuning, aligning ranking lists with LLM-preferred order is critical to enhance the final response quality. Therefore, we use an RL strategy to explore better ranking possibilities.
$\bullet$ Reward Function. We treat the quality of final responses as the reward of provided ranking lists to model LLM’s preferences. Since we expect the generated responses to cover all sub-answers from all sub-aspects, besides using the rouge score $\phi()$ to calculate the matching between responses $r$ and golden answers $a$, we further introduce a com-rouge score, $\phi^{c}()$ to measure the coverage of responses on sub-answers, $\mathcal{A}$. The reward function $\mathcal{r}(L)$ of a ranking list $L$ is produced as:
$\mathcal{r}(L)$ $=\phi(r,a)+\phi^{c}(r,\mathcal{A}),$ (11)
$\phi^{c}(r,\mathcal{A})$ $=\sum\nolimits_{i=1}^{n}\delta_{i}\phi(r,a_{i}),\;\;\;r=\mathcal{G}(q,L).$
$r$ is the response from the generator, $\mathcal{G}()$, given the query $q$ and ranked top-$k$ documents $L$. $\delta_{i}$ denotes the normalized length of the sub-answer to value the sub-answer’s weight.
Then, we adopt the Direct Preference Optimization (DPO) Rafailov et al. (2023) algorithm to ensure the optimization stability. Its training samples consist of a series of prediction pairs pre-generated by the policy model, namely the ranker in our study. Each pair contains a winner and a loser prediction, namely generated ranking lists, which are assessed by their rewards. Thus, DPO pairwise optimizes the policy model to discriminate the better one among a prediction pair.
$\bullet$ Data Construction. To build valuable training pairs, we introduce the unilateral sample significance strategy (US3). First, this approach generates a greedy search ranking list and multiple sampled ranking lists for each training data, obtaining their rewards via Equation (11). Then, US3 follows two rules when forming DPO training pairs: (1) Unilaterality: One prediction is from greedy search (used in inference) to provide a baseline for discerning better optimization directions, and the other from sampling search. (2) Significance: The reward gap between the predictions must exceed a threshold $\mu$ to ensure the pair’s value, thereby reducing errors from pairs with similar performance that may not reflect ranking quality, but noise.
$\bullet$ Optimization. Given built training pairs, we optimize the ranker using the following DPO objective function:
$\mathcal{L}_{DPO}=-\!\!\!\!\!\!\!\!\!\mathop{\mathbbm{E}}_{(x,y_{w},y_{l})\sim D}\!\!\!\!\!\![\log\sigma(\beta\log\frac{\pi_{\theta}(y_{w}|x){\pi_{f}(y_{l}|x)}}{\pi_{f}(y_{w}|x)\pi_{\theta}(y_{l}|x)}].$ (12)
$D$ is the training set built by US3, where the input of each data is $x=\{q,\hat{\mathcal{S}},\mathcal{P}\}$ and the output is a pair of ranking lists, with $y_{w}$ and $y_{l}$ as winner list and loser list respectively. $\pi_{\theta}$ denotes the policy model that needs to be optimized and $\pi_{f}$ represents the original policy model with fixed parameters, and its role is to avoid optimization trajectory excessively deviating from the basic model.
4 Experiment
4.1 Datasets and Metrics
Datasets. We conduct our experiments on two publicly available datasets that focus on multi-document summarization and long-form query-answer (QA) respectively, i.e., WikiPassageQA Hayashi et al. (2021) and WikiAsp Hayashi et al. (2021). WikiPassageQA offers human-annotated quality-evaluated questions and long-form answers. We chose this dataset because its answers are generally comprehensive, related to various aspects of questions, and the answer length is fairly long. WikiAsp dataset is devised for generating aspect-based summaries of entities from 20 domains. We follow Jiang et al. (2023) to convert it into open-domain QA settings. To support our experiments, we first operate some data pre-processing to ensure that each piece of data contains the question, ground truth answer, sub-aspects of the question, and their sub-answers. The process details and statistical information of datasets are demonstrated in Appendix A.
Metrics. To measure the matching scores of models’ responses with long-form ground truth answers, we select Rouge-1, Rouge-2, and Rouge-L as evaluation metrics. [^2] https://pypi.org/project/rouge/ Furthermore, we leverage com-rouge scores, which is introduced in Eq. 11, to assess the coverage of responses on sub-answers. We implement $\phi()$ in Eq. 11 by Rouge-1,-2, and -L to build Com-Rouge-1, -2, and -L evaluation metrics. For briefness, R1, R2, RL, CR1, CR2, and CRL are utilized to represent these metrics.
4.2 Baselines
To evaluate the effectiveness of our framework, we first build baselines with different RAG framework settings: (1) Close-book setting without external reference support. (2) “Retrieve-Generation” setting without ranking stage, namely No Ranker. (3) “Retrieve-Rerank-Generation” setting with various ranking algorithms, including RankT5 Nogueira et al. (2020); Zhuang et al. (2023a), a point-wise T5-based ranking model, LDist Izacard et al. (2024), a widely-used ranking algorithm in RAG systems Izacard et al. (2024); Shi et al. (2023), and LiT5 Tamber et al. (2023), a list-wise ranking model using slide-window-based ranking strategy. These baselines directly retrieve external documents based on the original queries. RAG-Fusion Rackauckas (2024) proposes retrieving documents from various sub-aspects and providing the final ranking lists via a simple reciprocal rank fusion algorithm Cormack et al. (2009). We set this framework as a basic baseline to compare the superiority of our proposed framework when explicitly considering query-aspects. We combine it with the above ranking algorithms to build various advanced versions of RAG-Fusion, e.g., RAG-Fusion+RankT5, etc. BGM Ke et al. (2024) is a recent RAG framework that introduces PPO strategy to fine-tune list-wise ranking model based on the LLM’s feedback. Due to limited space, we describe the implementation details in Appendix B.
4.3 Overall Results
To fully evaluate the effectiveness of our proposed framework, RichRAG, we utilize two settings to conduct experiments, the first one we provide the predicted sub-aspects to retriever and ranker while in the second one, we provide the golden sub-aspects to unlock the RichRAG’s powers in the fullest extent possible. We present the overall results in Table 1 and have the following conclusions:
Settings Models WikiPassageQA WikiAsp
R1 R2 RL CR1 CR2 CRL R1 R2 RL CR1 CR2 CRL
Predicted Sub-aspects Close-book .1350 .0324 .0989 .2052 .0397 .1714 .0913 .0114 .0581 .1189 .0128 .0834
No Ranker .2637 .1191 .2065 .3406 .1310 .2969 .1528 .0479 .0923 .1813 .0502 .1257
RankT5 .2719 .1313 .2123 .3486 .1441 .3027 .1529 .0481 .0932 .1813 .0504 .1263
LDIST .2729 .1325 .2140 .3497 .1447 .3044 .1528 .0479 .0923 .1996 .0624 .1376
LiT5 .2735 .1291 .2118 .3501 .1413 .3033 .1723 .0594 .1037 .2013 .0620 .1399
RAG-Fusion .2583 .1146 .2029 .3355 .1261 .2913 .1628 .0547 .0976 .1912 .0572 .1324
+RankT5 .2673 .1278 .2122 .3439 .1402 .3015 .1551 .0507 .0939 .1829 .0532 .1277
+LDIST .2669 .1253 .2114 .3445 .1394 .3022 .1711 .0587 .1021 .1996 .0612 .1380
+LiT5 .2649 .1226 .2071 .3418 .1347 .2987 .1696 .0589 .1017 .1978 .0616 .1373
BGM .2637 .1191 .2065 .3373 .1537 .2983 .1485 .0472 .0920 .1764 .0497 .1242
RichRAG .2930 .1538 .2316 .3688 .1664 .3225 .1839 .0678 .1094 .2127 .0706 .1458
Golden Sub-aspects No Ranker .3400 .1667 .2650 .4164 .1809 .3600 .1956 .0655 .1277 .2259 .0688 .1671
RankT5 .3535 .1760 .2734 .4270 .1887 .3673 .1923 .0632 .1244 .2228 .0665 .1637
LDIST .3543 .1766 .2728 .4296 .1906 .3680 .2144 .0792 .1354 .2441 .0826 .1767
LiT5 .3595 .1844 .2808 .4349 .1991 .3724 .2151 .0801 .1385 .2446 .0834 .1795
RAG-Fusion .3512 .1764 .2740 .4263 .1901 .3687 .2129 .0790 .1376 .2424 .0825 .1762
+RankT5 .3576 .1825 .2774 .4333 .1972 .3702 .2035 .0718 .1289 .2332 .0751 .1693
+LDIST .3616 .1851 .2787 .4358 .1996 .3734 .2164 .0791 .1366 .2459 .0825 .1767
+LiT5 .3646 .1906 .2842 .4393 .2053 .3776 .2165 .0800 .1369 .2457 .0833 .1770
BGM .3393 .1686 .2723 .4161 .1824 .3651 .1922 .0642 .1282 .2231 .0677 .1666
RichRAG .3973 .2247 .3055 .4688 .2392 .4015 .2321 .0942 .1456 .2606 .0976 .1863
Table 1: Overall results of all models. The best and second-best results are in bold and underlined, respectively.
Models WikiPassageQA
R1 R2 RL CR1 CR2 CRL
RichRAG .3973 .2247 .3055 .4688 .2392 .4015
w/o SA .3262 .1505 .2610 .4046 .1648 .3516
w/o GLR .3467 .1706 .2703 .4229 .1850 .3638
w/o RL .3854 .2100 .2946 .4584 .2248 .3907
w/o US3 .3853 .2102 .2954 .4583 .2251 .3910
Models WikiAsp
R1 R2 RL CR1 CR2 CRL
RichRAG .2321 .0942 .1456 .2606 .0976 .1863
w/o SA .2096 .0678 .1479 .2058 .0541 .1557
w/o GLR .1742 .0504 .1190 .2441 .0806 .1737
w/o RL .2275 .0884 .1422 .2566 .0919 .1824
w/o US3 .2189 .0833 .1381 .2482 .0867 .1773
Table 2: Ablation results of RichRAG on two datasets.
(1) Whether given predicted or golden sub-aspects, RichRAG shows the best performance. This phenomenon confirms the ability of our framework to explore and leverage user’s sub-intents underlying the issued multi-faceted questions, hence providing comprehensive responses. However, existing RAG systems solely consider query-document relevance without relationships among candidates, blocking their potential to understand user’s various sub-intents and limiting the richness of final responses.
(2) Compared to list-wise ranking algorithms, our method still illustrate better performance. Though there exist several list-wise ranking algorithms in the RAG community, such as LiT5 and BGM, these algorithms do not explicitly model the interactions among candidates from the perspective of comprehensiveness of the user intent coverage. Without such explicit guidance, these algorithms may be trapped in a locally optimal solution, hence impacting the overall quality of ranking lists.
(3) With golden sub-aspects, RAG-Fusion settings often outperform settings only considering the original question’s retrieved documents. It reveals the importance of modeling the questions’ sub-aspects in RAG systems for generating rich and reliable responses. However, with predicted sub-aspects, the RAG-Fusion variants do not outperform corresponding baselines without RAG-Fusion. This may be due to the gap between predicted sub-aspects and annotated sub-aspects. However, in real applications, the user’s sub-intents may be diverse while we can only label some of them in datasets to evaluate model performances. Therefore, how to deal with such a gap between realistic and human annotation is still an open problem and will be further investigated in our future study.
4.4 Ablation Studies
In order to evaluate the role of our key modules, we further conduct the following ablation studies with results presented in Table 2.
(1) To confirm the importance of explicitly consideration of user’s sub-intents, i.e.question’s sub-aspects. We further construct a variant, w/o SA, by directly ranking the retrieved documents of the original question without considering the candidate pool, $\mathcal{P}$. The significantly worse results than RichRAG further prove the importance of explicitly considering the various sub-intents underlying multi-faceted questions, which is beneficial for providing comprehensive responses for users.
(2) In our study, we propose a generative list-wise ranking module to generate LLM-preferred comprehensive ranking lists. To prove its advantages, we replace it with another list-wise algorithm, LiT5, in our framework to build a variant, w/o GLR. We find that it still underperforms RichRAG. This result validates the advantages of our model structure, i.e.the ability to potentially model global interactions among various candidates with sub-aspects. While the sliding-window-based list-wise algorithm still has defects on it, hence limiting the performance.
(3) To confirm the role of alignment with LLMs’ preferences, we build a variant of our framework, w/o. RL, which only supervised fine-tunes our ranker without the RL optimization. The declined performance proves that the LLMs’ preferences are different from humans’ preferences. As a result, it is vital for RAG systems to align the LLMs’ preferences to enhance the overall quality of the final responses generated from LLMs.
(4) To ensure the robustness and quality of the DPO algorithm, we propose a US3 approach to build the pairwise training samples for it. To confirm its effect, we replace it by randomly creating the training pairs for the RL stage, building a variant, namely w/o US3. The worse result than RichRAG proves the usefulness of this strategy. It suggests that the US3 approach can create more reliable training pairs by ensuring the meaningful comparison between predictions of our ranker, hence optimizing it in a promising direction.
Figure 3: Subset experiments with different sub-aspect amounts.
4.5 Impact of Sub-aspect Amount
To test the generalization of our framework with different sub-aspects numbers, which represent different search scenarios, we further split the test dataset into different sub-sets according to the number of sub-aspects. Questions with a sub-aspect amount less than two are divided into the narrow set, questions with a sub-aspect amount less than four are divided into the middle set, and the remaining questions are divided into the broad set. The models’ performances on these subsets are shown in Figure 3. Evidently, our framework outperforms all baselines on all subsets. This result verifies the robustness of our framework with diverse search scenarios. Furthermore, we find that the overall results on the broad set are worse than the remaining two sets. This phenomenon implies that the scenarios involving various potential user sub-intents are harder to handle than specific user intents, which need to be further investigated in the future.
4.6 Efficiency Analysis of Ranking Algorithms
We previously confirmed the advantages of considering various sub-aspects in RAG systems. However, with extensive candidate documents, the efficiency of ranking modules is also important. Therefore, we compare the query latency of our ranker to point-wise and list-wise ranking algorithms, LDIST and LiT5, to test their efficiency. First, we demonstrate their changes in query latency with the candidate amount in Figure 4 (a). Obviously, our ranker has comparable efficiency and trend with the point-wise ranking algorithm. However, the time overhead of LiT5 rises more sharply along with the improvement of the candidate amount. This phenomenon proves that our ranker could provide a better trade-off between effectiveness and efficiency. Furthermore, since our ranker generates docids step-by-step, we further provide the trends of query latency with different generated document numbers and show the trend lines of different amounts of candidate documents in Figure 4 (b). It can be found that as the number of ranked documents increases, all trendlines rise slowly, and the gap between different candidate counts (CCnt) is limited to 1 second. This phenomenon further proves the robustness of our ranker’s efficiency across diverse ranking settings.
Due to limited space, we provide further analysis studies in Appendix C,  D and  E.
Figure 4: Efficiency experiments of different models.
5 Conclusion
In this study, we proposed a new RAG framework, RichRAG, to comprehensively consider the various sub-intents underlying users’ broad questions, hence providing all-sided long-form responses for users. Specifically, we introduced a sub-aspect explorer to predict the potential sub-aspects contained by questions representing the user’s sub-intents. According to sub-aspects and a fixed retriever, we could build extensive and diverse candidate pools. To provide comprehensive and LLM-preferred ranking lists, we designed a generative list-wise ranking model. It effectively and efficiently encodes the global relationships between candidates and multi-aspects, thereby offering global optimal ranking lists to LLMs. To ensure the ranking quality, we utilized a two-stage training process involving supervised fine-tuning and RL optimization. Furthermore, we devised a US3 approach to create useful and reliable training samples to ensure the effectiveness of the DPO algorithm. Extensive experiments on two public datasets confirm the effectiveness and efficiency of RichRAG.
Limitations
In this work, we identified an underexplored but important scenario of RAG systems, where multi-faceted questions require rich and comprehensive responses satisfying various related sub-aspects. To handle these situations, we developed a framework, namely RichRAG to equip RAG models with the ability to generate rich and satisfying responses for multi-faceted questions. We acknowledge the following limitations of our current study that present opportunities for future investigations.
First, though we built an aspect explorer to identify users’ sub-intents underlying multi-faceted questions, it is still shallow and there is a gap between predicted sub-aspects and real intents. This is because the user’s intents are usually diverse and vary from person to person. Even though we annotated some sub-aspects in data samples, these may only cover a sub-set. Therefore, in this study, we mainly focus on how to provide a promising reference permutation given the user’s potential sub-intents for enhancing the final generation to comprehensively respond to these sub-intents. Second, since the situation is still underexplored, few suitable datasets are available. The datasets we used in our study were chosen by carefully investigating the data samples’ content and converted by some operations to a suitable data format. Therefore, the diversity of experiment datasets is limited. In the future, we will pay more attention to the evaluation and annotation of user intent exploration in such scenarios to support further study.
Appendix A Data Pre-process
Specifically, for the WikiPassageQA dataset, we prompt GPT-4 to generate appropriate and accurate sub-aspects and sub-answers for each data point to support our study, where sub-answers are split from the original long-form answers and sub-aspects are closely related to original questions and sub-answers. The content of the prompt is shown in Prompt E. The WikiAsp dataset is devised for generating aspect-based summaries of entities from 20 domains. Each piece of data is built from a Wikipedia article, consisting of various aspects of this article and aspect-based summaries, each target summary corresponds to an aspect. We follow Jiang et al. (2023) to convert it into open-domain QA settings and introduce some additional operations to make it suitable for our experimental settings. Firstly, samples with more than two aspects are retained and the remaining parts are removed to ensure the multifaceted nature of the experimental samples. Then, due to the expensive costs of experiments and the large amount of the whole dataset, we evenly sample subsets from each domain to build the experimental samples, and split training, validation, and test sets according to the ratio of 10:1:1. Finally, we insert the title of the original Wikipedia article into a template: “Generate a summary about title” to mimic the real question format, hence constructing the question of each sample data. The aspects and aspect-based summaries are treated as sub-aspects of the question and sub-answers. We concatenate these sub-answers to build the long-form answer for each sample. The statistical information of datasets is presented in Table 3.
Items WikiPassageQA
Train Validation Test
Count 3,311 415 416
Avg. Q Len 9.53 9.70 9.44
Avg. A Len 148.14 145.93 146.1
Avg. SubCnt 3.77 3.77 3.78
Avg. Sub Q Len 6.34 6.25 6.29
Avg. Sub A Len 62.84 62.32 61.99
Items WikiAsp
Train Validation Test
Count 8,613 859 867
Avg. Q Len 7.01 6.97 6.94
Avg. A Len 201.7 216.26 200.44
Avg. SubCnt 2.38 2.41 2.40
Avg. Sub Q Len 1.28 1.28 1.29
Avg. Sub A Len 229.36 240.6 221.91
Table 3: Statistical information of datasets.
Appendix B Implementation Details
The sub-aspect explorer is implemented by Llama-2-7B-chat Touvron et al. (2023). We set the learning rate as 5e-5, batch size as 64, and use AdamW Loshchilov and Hutter (2019) to fine-tune it. For the multi-faceted retriever, the number of retrieved documents for each sub-aspect is set as 50. The maximum capacity of the pool is 290 for WikipasssageQA and 270 for WikiAsp. For the generative list-wise ranker, We base on Flan-T5-base Chung et al. (2022) to initialize it and rerank the top-10 final documents as provided external knowledge for the generator. In the SFT stage, we set the learning rate as 5e-5, batch size as 64, temperature $\tau$ as 0.1, and optimize the ranker to generate top-$10$ ranked document IDs with AdamW algorithm. In the RL stage, Llama-2-13B-chat is chosen as the generator, $\mathcal{G}$, providing reward feedback to the policy model. Then, we set $\mu$ as 0.1 to build 6,000 training pairs for the DPO algorithm. The batch size and learning rate are set as 32 and 3e-6 respectively to further optimize our ranker via the DPO strategy. We followed Izacard et al. (2024) to consider the Dec. 20, 2021, Wikipedia dump as our knowledge base and utilize BGE-en-base Xiao et al. (2023) as our fixed retriever. Our experiments are conducted on the platform with four NVIDIA A100-SXM4-80GB GPUs. We will release our codes upon the acceptance of our study.
Figure 5: Trend of model performance as Top-K changes on both two dataset.
Appendix C Analysis of LLMs’ Preferences
We mentioned that in the system of RAG, the downstream users of IR systems are no longer humans, but LLMs. To align the LLMs’ reading preferences on the provided information, we introduce the RL stage to further capture LLMs’ preferences on the order of ranking lists. Furthermore, another angle of the differences between human and LLM users is that traditional IR systems usually provide distinct documents for users and assume that users will carefully read the document containing information if she is interested in a certain document. However, such a reading habit may not be suitable for LLMs. It may be important for LLMs to repeat some important information when providing retrieved knowledge Ke et al. (2024). Such a paradigm is hard to implement by traditional ranking models based on individual relevance score sorting. However, it is easy to accomplish for our generative ranking model. Therefore, we waive the constraint of ensuring that the next ranked document has not been previously selected. The corresponding results are illustrated in Figure 6. Interestingly, releasing of repetition constraint could bring significant improvement to our model. The potential reason may be that repetition of important information could avoid the introduction of irrelevant information and attract more attention of LLMs to repeated important information, which enhances the LLMs’ confidence in it. This enables LLMs to provide more reliable responses according to this important knowledge. Similar results can also be found in Wang et al. (2024b). It further confirms the importance of repetition with potential emphasis and denoising effect.
Appendix D Impact of Number of External Documents
To further investigate the impact of different numbers of external knowledge on RAG performance, we vary the value of K and conduct corresponding experiments on our model and a baseline model that directly treats the top-k retrieved documents as provided references for the generator. We set the maximum value to 14 due to the limited input length of the generator. The performance trendlines of the two models are shown in Figure 5. We further provide the trendlines of our model’s improvement at each point with red lines. By comparison, we find that our model generally outperforms the baseline with different top-k numbers, which confirms the superiority of RichRAG. In addition, it is clear that the performance of the baseline usually improves with the initial increase in top-k values. This phenomenon suggests that the baseline cannot rank the valuable documents at the forefront. Therefore, important documents can only be incorporated into the generator’s input when the top value becomes larger. However, our model is capable of accurately ranking the valuable documents at the top of the ranking. Consequently, even with a small k-value, it still demonstrates excellent performance. The improvement trendlines also imply that with limited external references, our model could show better response performance due to better ranking abilities.
Appendix E Case Study
To validate that our proposed framework has the ability to provide rich and comprehensive responses for multi-faceted questions, we further demonstrate a case study to compare the generated responses of RichRAG and a RAG baseline that directly generates responses based on retrieved documents. In Box E, we provide the input question, its labeled sub-aspects, and predictions from different compared models. The parts of generated responses are highlighted by different colors to indicate the comprehensiveness of different results. According to the visualized results, it is obvious that RichRAG could predict the query’s sub-aspects accurately and provide rich responses that satisfy these potential intents. However, the baseline model only offers a general and incomplete response without in-depth consideration of the various underlying intentions of the user, which may implicitly impact user satisfaction. This result further proves the ability of RichRAG to identify and explicitly model the query’s sub-aspects, thereby providing satisfying ranking results to enhance the richness and comprehensiveness of generated responses.
Figure 6: Comparison with and without repeated constraints
Box E: A case to show the richness and comprehensiveness of our model’s generated responses. Query: Generate a summary about Instant Mom.
Sub-aspects: production, reception
Baseline Response:
Instant Mom is a talk show that focuses on various aspects of motherhood, including child-rearing, homemaking, and personal growth. The show features a variety of segments, including MomStop, which provides pressure-free bonding moments between mothers and their children, and Momstyle, which showcases different techniques in child rearing and domestic engineering. The show also includes 24/7, which celebrates the non-stop nature of motherhood, and Mom and Me, an inspiring tribute to the kindness and love of mothers. Additionally, the show features practical tips for homemakers and parents, as well as interviews with celebrity mothers and experts in child development. The show is certified by moms and features independent reviews and expert advice to provide a comprehensive resource for mothers.
RichRAG Response:
Based on the given sources, here is a summary of Instant Mom:
Instant Mom is a television show that aired on Nickelodeon from 2013 to 2015. The show was created by Howard Nemetz and starred Tia Mowry-Hardrict as Stephanie, a young woman who becomes the legal guardian of her three nieces and nephews after her sister and brother-in-law die in a car accident. The show follows the challenges and adventures of Stephanie and her new family as they navigate life and try to move forward after their loss. The show received positive reviews, with Emily Ashby of Common Sense Media giving it 4 out of 5 stars and Brian Lowry of Variety comparing it to Trophy Wife. The show also featured a production report, which is a filmmaking term for a form filled out each day of production to summarize what occurred. The show was produced by Nickelodeon and was filmed in various locations in Los Angeles. Overall, Instant Mom is a heartwarming and humorous show that explores the complexities of family and the challenges of raising children. With its talented cast and engaging storylines, it is a great choice for families and young audiences looking for a fun and entertaining show.
Prompt E: Annotate question aspects and split long-form answers into corresponding sub-answers Your task is to adjust the results of query-facets mining. The query-facets are extensions of the original query in various generic perspectives, rather than some specific facts. Given a query that requires information from multiple query-facets, you should return all query-facets of the query to fully answer it query. Note that each query has at least two query-facets. I will give you the long-form answer to the original query to help you explore query-facets based on the perspectives of its answer. But refrain from using the additional information from the answer to generate the query-facets. Then you should segment the original long-form answer into several sub-answers that each are paired with a query-facet. Please return each query-facet of the original query and its corresponding sub-answers. The query-facets and sub-answers should be one-to-one and returned in JSON format. You need to follow the following rules:
1. The answers are only used to help you determine the generic direction. You mustn’t generate query-facets based on the contents of answers and the facets mustn’t contain the answers’ additional information beyond the input query.
2. Sub-answers are constructed by segmenting the original answer, you cannot generate or reorder the original answer to create sub-answers.
3. The sub-answers should be complete. You must ensure that when the sub-answers are joined together in order, the complete original answer should be formed.
4. The generated query-facets should be sufficiently generic and contain no specific information about the sub-answers.
5. **You should ensure that generated query-facets cover all perspectives original answer.**
6. **You should ensure that all sub-answers cover all contents of the original answer.**
7. **The number of query surfaces must range from 2 to 7.**
8. **You should ensure that each query-facet is sufficiently generic and can be easily derived from the original query.**
9. **You should ensure each query-facet contains no information from the answer.**
10. **You should rewrite or combine the query-facets to be more generic if some query-facets do not meet the above requirements.**
11. The returned results should be in JSON format and contain the following key: results, which is a list of JSON data. Each item of results should contain the following keys: query-facet, and sub-answer.
12. I will give you some demonstrations, you should learn the pattern of them to mine query-facets and split sub-answers.
**Demonstration**
{demonstrations}
Query: {query}
Answer: {answer}
Results:

### v3

@@
 Shuting Wang1, Xin Yu2, Mang Wang2, Weipeng Chen2, Yutao Zhu1and Zhicheng Dou1∗ 1Gaoling School of Artificial Intelligence, Renmin University of China 2Baichuan Intelligent Technology {wangshuting, dou}@ruc.edu.cn
-**footnotetext: Corresponding author.
 1 Introduction
@@
 (2024); Wang et al.
-(2023c); Li et al.
+(2023d); Li et al.
 (2024) have been proposed, existing studies primarily focus on addressing specific problems that require concise and definitive answers.
@@
 However, such redundant candidates inevitably contain much irrelevant noise and are hard to handle completely by LLMs due to limited input length.
-As a result, sorting out the top-k most valuable documents from the candidate pool is critical to the success of the RichRAG framework.
+As a result, sorting out the top-k best documents from the candidate pool is critical to the success of the RichRAG framework.
 In further, we claim that a promising top-k ranking should have the following desirable features: (1) Comprehensiveness.
 Incentivizing LLM to generate rich and reliable responses requires external documents to comprehensively cover various query aspects.
-Therefore, the ability to model relationships among documents, hence maximizing the coverage of the entire reference list is crucial for the ranking module.
+Therefore, instead of predicting each document’s relevance independently, the ranking model should consider the relationship among documents to enhance the global coverage of the ranking list for query aspects.
 (2) Alignment with the LLMs’ preferences.
@@
 This structure offers two key advantages: (1) Global Document Modeling.
-The seq-to-seq model structure equips the ranker to effectively model global interactions among candidate documents, queries, and sub-aspects, thereby capturing the overall utility of generated ranking lists in covering the query’s multi-aspects.
+The seq-to-seq model structure equips the ranker to effectively model global interactions among candidates, queries, and sub-aspects, thereby capturing the overall utility of generated ranking lists in covering the query’s multi-aspects.
 (2) Efficiency.
@@
 The optimization of our ranker consists of two stages: The first is supervised fine-tuning (SFT).
-To enhance the coverage of generated ranking lists on query aspects, we devise a coverage utility function based on which to build silver generation targets for training samples greedily.
+To enhance the coverage of generated ranking lists on query aspects, we devise a coverage utility function based on which to build silver generation targets (ranking lists of document IDs) for training samples greedily.
 These silver targets allow us to SFT our ranker and ensure its basic ability.
 To further improve the ranking quality and align ranking lists with LLMs’ preferences, a reinforcement learning stage is introduced.
-We measure the LLMs’ final response quality to create reward values and adopt the DPO Rafailov et al.
+We consider both the accuracy and comprehensiveness of generated responses to create reward values, and adopt the DPO Rafailov et al.
 (2023) algorithm to optimize our ranker.
@@
 2.1 Retrieval-augmented Generation
-The RAG paradigm has exhibited excellent abilities in enhancing LLMs to generate accurate responses.
-To effectively optimize the system, some studies focus on concurrently or asynchronously training retrievers and generators Izacard et al.
+To ensure the effectiveness of RAG systems, previous studies mainly optimized retrievers and generators simultaneously Izacard et al.
 (2024); Borgeaud et al.
@@
 (2024).
-However, as LLMs grow in size and inference costs rise, researchers have also explored fixing LLMs and optimizing retrievers as plug-in modules Shi et al.
+Recent researchers also explored fixing LLMs and optimizing retrievers as plug-in modules Shi et al.
 (2023); Yu et al.
@@
 (2023); Wang et al.
-(2023c); Li et al.
+(2023d); Li et al.
 (2024); Ke et al.
 (2024); Gao et al.
-(2024).
+(2024), to reduce the training cost.
 Some studies Chan et al.
@@
 (2023); Yao et al.
-(2023) propose to decompose multi-hop questions into sub-tasks and solve them step-by-step to produce specific answers.
-However, our study mainly focuses on the query scenarios with various implicit sub-intents that require rich and comprehensive responses.
-It is different from the query decomposition scenario.
+(2023); Xu et al.
+(2024a) propose to decompose multi-hop questions into sub-tasks and solve them step-by-step.
+These works typically focus on breaking down complex questions with clearly stated user intents into simpler questions.
+Our research, however, addresses a different scenario where user questions are broad and encompass various potential sub-aspects not explicitly stated.
+Answering such questions usually requires integrating diverse relevant information to these underlying sub-aspects to fully respond to the user’s potential information needs.
 Recent studies Rackauckas (2024) also highlight the importance of exploring users’ sub-intents, but its simplistic pipeline fails to model global document-intent interactions, leading to sub-optimal results.
@@
 (2024); Xu et al.
-(2024) also explore list-wise rankers in RAG systems but still focus on scenarios with specific intents and answers, neglecting the depth and breadth of user questions.
+(2024b) also explore list-wise rankers in RAG systems but still focus on scenarios with specific intents and answers, neglecting the depth and breadth of user questions.
 3 Method
@@
 We leverage LLMs to build our sub-aspect explorer, $\mathcal{E}$, due to their extensive world knowledge and excellent capabilities in language understanding and generation.
-This module takes a prompt $p_{se}$, which instructs the LLM to predict the sub-aspects of the input query, and a user’s query, $q$, as input and generates a series of sub-aspects under the query, i.e., ${\hat{\mathcal{S}}}=\{\hat{s}_{1},\ldots,\hat{s}_{m}\}$, as follows:
-$\{\hat{s}_{1},\ldots,\hat{s}_{m}\}=\mathcal{E}(q,p_{se}).$ (1)
+This module takes a prompt $p_{se}$, which instructs the LLM to predict the sub-aspects of the input query, and a user’s query, $q$, as input and generates a series of sub-aspects under the query:
+$\vspace{-2pt}{\hat{\mathcal{S}}}=\{\hat{s}_{1},\ldots,\hat{s}_{m}\}=\mathcal{E}(q,p_{se}).$ (1)
 To align the sub-aspect explorer with the output format and the distribution of downstream data, we fine-tune it using training queries and their labeled sub-aspects.
@@
 To equip our ranker with the ability to globally model relationships among candidates, we build it upon a generative model, T5 Raffel et al.
-(2020), which views all candidates, sub-aspects, and the query as input and directly generates a top-k ranking list of document IDs (docids).
+(2020), which views all candidates, sub-aspects, and the query as input and directly generates a top-k ranking list of document ID (docid) tokens.
 For each candidate, $d_{i}$, we concatenate it with the original query, $q$, its associated sub-aspects, $s(d_{i})=\{\hat{s}^{i}_{1},\ldots,\hat{s}^{i}_{n_{i}}\}$, and some special tokens to formulate an input sequence:
@@
 $\mathrm{\mathbf{e}}_{i}$ $=\mathrm{Pool}(\mathrm{Enc}(I_{i})),$
-where $r(i)$ project the rank position $i$ into the generated docid that is ranked at the $i$-th position. This pooling operation could significantly reduce the time-space burden of the decoder.
+where $r(i)$ project the rank position $i$ into the index of the document ranked at the $i$-th position. This pooling operation could significantly reduce the time-space burden of the decoder.
 Additionally, we implement a reuse strategy on the language model (LM) head layer to reduce unnecessary load and enhance the modeling accuracy.
@@
 $w^{t}_{i}=1-\mathrm{Norm}_{i}({\max\nolimits_{\tilde{d}\in L^{*}_{t-1}}\phi(\tilde{d},a_{i})}),$ (9)
-The silver target list, $L^{*}_{k}$, allows us to supervised fine-tune the ranker via the NTP task:
+The silver target list, $L^{*}_{k}$, allows us to supervise fine-tune the ranker via the NTP task:
 $\mathcal{L}_{sft}=-\sum\nolimits_{t=1}^{k}\log p(d^{*}_{t}|q,\mathcal{P},\hat{\mathcal{S}},L^{*}_{t-1}).$ (10)
+$p(d^{*}_{t}|q,\mathcal{P},\hat{\mathcal{S}},d^{*}_{1:t-1})$ is the generation probability of the $t$-th target docid conditioned on the current question, the candidate pool, the sub-aspects, and the previously target documents.
+It is calculated by our ranking module.
 3.4.2 Reinforcement Learning
@@
 $\pi_{\theta}$ denotes the policy model that needs to be optimized and $\pi_{f}$ represents the original policy model with fixed parameters, and its role is to avoid optimization trajectory excessively deviating from the basic model.
+Settings Models WikiPassageQA WikiAsp
+F1 R2 RL BS CR2 CRL F1 R2 RL BS CR2 CRL
+Predicted Sub-aspects Close-book .2132 .0324 .0989 .8417 .0397 .1714 .2037 .0114 .0581 .7967 .0128 .0834
+No Ranker .3372 .1191 .2065 .8491 .1310 .2969 0.3029 .0479 .0923 .8073 .0502 .1257
+RankT5 .3502 .1313 .2123 .8536 .1441 .3027 0.3079 .0481 .0932 .8063 .0504 .1263
+LDIST .3487 .1325 .2140 .8523 .1447 .3044 .3377 .0479 .0923 .8114 .0624 .1376
+LiT5 .3473 .1291 .2118 .8514 .1413 .3033 .3341 .0594 .1037 .8125 .0620 .1399
+RAG-Fusion .3316 .1146 .2029 .8467 .1261 .2913 .3219 .0547 .0976 .8116 .0572 .1324
++RankT5 .3448 .1278 .2122 .8510 .1402 .3015 .3123 .0507 .0939 .8091 .0532 .1277
++LDIST .3400 .1253 .2114 .8494 .1394 .3022 .3327 .0587 .1021 .8122 .0612 .1380
++LiT5 .3386 .1226 .2071 .8491 .1347 .2987 .3344 .0589 .1017 .8137 .0616 .1373
+BGM .3465 .1191 .2065 .8547 .1537 .2983 .2969 .0472 .0920 .8061 .0497 .1242
+RichRAG .3638 .1538 .2316 .8549 .1664 .3225 .3610 .0678 .1094 .8194 .0706 .1458
+Golden Sub-aspects No Ranker .3763 .1667 .2650 .8544 .1809 .3600 .3325 .0655 .1277 .8116 .0688 .1671
+RankT5 .3854 .1760 .2734 .8569 .1887 .3673 .3329 .0632 .1244 .8107 .0665 .1637
+LDIST .3867 .1766 .2728 .8558 .1906 .3680 .3609 .0792 .1354 .8163 .0826 .1767
+LiT5 .3926 .1844 .2808 .8558 .1991 .3724 .3564 .0801 .1385 .8165 .0834 .1795
+RAG-Fusion .3808 .1764 .2740 .8554 .1901 .3687 .3658 .0790 .1376 .8187 .0825 .1762
++RankT5 .3866 .1825 .2774 .8572 .1972 .3702 .3511 .0718 .1289 .8155 .0751 .1693
++LDIST .3894 .1851 .2787 .8610 .1996 .3734 .3690 .0791 .1366 .8190 .0825 .1767
++LiT5 .3915 .1906 .2842 .8567 .2053 .3776 .3676 .0800 .1369 .8190 .0833 .1770
+BGM .3723 .1686 .2723 .8537 .1824 .3651 .3296 .0642 .1282 .8104 .0677 .1666
+RichRAG .4174 .2247 .3055 .8637 .2392 .4015 .3951 .0942 .1456 .8256 .0976 .1863
+Table 1: Overall results of all models. The best and second-best results are in bold and underlined, respectively.
 4 Experiment
@@
 Metrics.
-To measure the matching scores of models’ responses with long-form ground truth answers, we select Rouge-1, Rouge-2, and Rouge-L as evaluation metrics.
-[^2] https://pypi.org/project/rouge/ Furthermore, we leverage com-rouge scores, which is introduced in Eq.
+To measure the matching scores of models’ responses with long-form ground truth answers, we select F1, Rouge(-2 and -L), and BERT-Score as evaluation metrics.Furthermore, we leverage the com-rouge score, which is introduced in Eq.
 11, to assess the coverage of responses on sub-answers.
 We implement $\phi()$ in Eq.
-11 by Rouge-1,-2, and -L to build Com-Rouge-1, -2, and -L evaluation metrics.
-For briefness, R1, R2, RL, CR1, CR2, and CRL are utilized to represent these metrics.
+11 by Rouge-2, and -L to build Com-Rouge-2 and -L evaluation metrics.
+For briefness, F1, R2, RL, BS, CR2, and CRL are utilized to represent these metrics.
+Following Wang et al.
+(2023c), we also leverage GPT-4 to conduct a pairwise evaluation of our method with baseline models to confirm its effectiveness further.
+To comprehensively evaluate the effectiveness of our proposed ranking algorithm, we also provide the ranking performance comparison and analysis in Appendix C
+Figure 3: Results of the GPT-4-based evaluation comparing our method with baseline models.
 4.2 Baselines
@@
 To fully evaluate the effectiveness of our proposed framework, RichRAG, we utilize two settings to conduct experiments, the first one we provide the predicted sub-aspects to retriever and ranker while in the second one, we provide the golden sub-aspects to unlock the RichRAG’s powers in the fullest extent possible.
-We present the overall results in Table 1 and have the following conclusions:
-Settings Models WikiPassageQA WikiAsp
-R1 R2 RL CR1 CR2 CRL R1 R2 RL CR1 CR2 CRL
-Predicted Sub-aspects Close-book .1350 .0324 .0989 .2052 .0397 .1714 .0913 .0114 .0581 .1189 .0128 .0834
-No Ranker .2637 .1191 .2065 .3406 .1310 .2969 .1528 .0479 .0923 .1813 .0502 .1257
-RankT5 .2719 .1313 .2123 .3486 .1441 .3027 .1529 .0481 .0932 .1813 .0504 .1263
-LDIST .2729 .1325 .2140 .3497 .1447 .3044 .1528 .0479 .0923 .1996 .0624 .1376
-LiT5 .2735 .1291 .2118 .3501 .1413 .3033 .1723 .0594 .1037 .2013 .0620 .1399
-RAG-Fusion .2583 .1146 .2029 .3355 .1261 .2913 .1628 .0547 .0976 .1912 .0572 .1324
-+RankT5 .2673 .1278 .2122 .3439 .1402 .3015 .1551 .0507 .0939 .1829 .0532 .1277
-+LDIST .2669 .1253 .2114 .3445 .1394 .3022 .1711 .0587 .1021 .1996 .0612 .1380
-+LiT5 .2649 .1226 .2071 .3418 .1347 .2987 .1696 .0589 .1017 .1978 .0616 .1373
-BGM .2637 .1191 .2065 .3373 .1537 .2983 .1485 .0472 .0920 .1764 .0497 .1242
-RichRAG .2930 .1538 .2316 .3688 .1664 .3225 .1839 .0678 .1094 .2127 .0706 .1458
-Golden Sub-aspects No Ranker .3400 .1667 .2650 .4164 .1809 .3600 .1956 .0655 .1277 .2259 .0688 .1671
-RankT5 .3535 .1760 .2734 .4270 .1887 .3673 .1923 .0632 .1244 .2228 .0665 .1637
-LDIST .3543 .1766 .2728 .4296 .1906 .3680 .2144 .0792 .1354 .2441 .0826 .1767
-LiT5 .3595 .1844 .2808 .4349 .1991 .3724 .2151 .0801 .1385 .2446 .0834 .1795
-RAG-Fusion .3512 .1764 .2740 .4263 .1901 .3687 .2129 .0790 .1376 .2424 .0825 .1762
-+RankT5 .3576 .1825 .2774 .4333 .1972 .3702 .2035 .0718 .1289 .2332 .0751 .1693
-+LDIST .3616 .1851 .2787 .4358 .1996 .3734 .2164 .0791 .1366 .2459 .0825 .1767
-+LiT5 .3646 .1906 .2842 .4393 .2053 .3776 .2165 .0800 .1369 .2457 .0833 .1770
-BGM .3393 .1686 .2723 .4161 .1824 .3651 .1922 .0642 .1282 .2231 .0677 .1666
-RichRAG .3973 .2247 .3055 .4688 .2392 .4015 .2321 .0942 .1456 .2606 .0976 .1863
-Table 1: Overall results of all models. The best and second-best results are in bold and underlined, respectively.
-Models WikiPassageQA
-R1 R2 RL CR1 CR2 CRL
-RichRAG .3973 .2247 .3055 .4688 .2392 .4015
-w/o SA .3262 .1505 .2610 .4046 .1648 .3516
-w/o GLR .3467 .1706 .2703 .4229 .1850 .3638
-w/o RL .3854 .2100 .2946 .4584 .2248 .3907
-w/o US3 .3853 .2102 .2954 .4583 .2251 .3910
-Models WikiAsp
-R1 R2 RL CR1 CR2 CRL
-RichRAG .2321 .0942 .1456 .2606 .0976 .1863
-w/o SA .2096 .0678 .1479 .2058 .0541 .1557
-w/o GLR .1742 .0504 .1190 .2441 .0806 .1737
-w/o RL .2275 .0884 .1422 .2566 .0919 .1824
-w/o US3 .2189 .0833 .1381 .2482 .0867 .1773
-Table 2: Ablation results of RichRAG on two datasets.
+We present the overall results in Table 1 and Figure 3, and have the following conclusions:
 (1) Whether given predicted or golden sub-aspects, RichRAG shows the best performance.
@@
 However, existing RAG systems solely consider query-document relevance without relationships among candidates, blocking their potential to understand user’s various sub-intents and limiting the richness of final responses.
-(2) Compared to list-wise ranking algorithms, our method still illustrate better performance.
+(2) Compared to list-wise ranking algorithms, our method still illustrates better performance.
 Though there exist several list-wise ranking algorithms in the RAG community, such as LiT5 and BGM, these algorithms do not explicitly model the interactions among candidates from the perspective of comprehensiveness of the user intent coverage.
@@
 4.4 Ablation Studies
-In order to evaluate the role of our key modules, we further conduct the following ablation studies with results presented in Table 2.
+In order to evaluate the role of our key modules, we further conduct the following ablation studies with results presented in Figure 4.
+This figure shows the decline degrees of the ablation models compared with the complete RichRAG.
 (1) To confirm the importance of explicitly consideration of user’s sub-intents, i.e.question’s sub-aspects.
@@
 The significantly worse results than RichRAG further prove the importance of explicitly considering the various sub-intents underlying multi-faceted questions, which is beneficial for providing comprehensive responses for users.
+Figure 4: Ablation results of RichRAG on two datasets.
 (2) In our study, we propose a generative list-wise ranking module to generate LLM-preferred comprehensive ranking lists.
@@
 It suggests that the US3 approach can create more reliable training pairs by ensuring the meaningful comparison between predictions of our ranker, hence optimizing it in a promising direction.
-Figure 3: Subset experiments with different sub-aspect amounts.
-4.5 Impact of Sub-aspect Amount
-To test the generalization of our framework with different sub-aspects numbers, which represent different search scenarios, we further split the test dataset into different sub-sets according to the number of sub-aspects.
-Questions with a sub-aspect amount less than two are divided into the narrow set, questions with a sub-aspect amount less than four are divided into the middle set, and the remaining questions are divided into the broad set.
-The models’ performances on these subsets are shown in Figure 3.
-Evidently, our framework outperforms all baselines on all subsets.
-This result verifies the robustness of our framework with diverse search scenarios.
-Furthermore, we find that the overall results on the broad set are worse than the remaining two sets.
-This phenomenon implies that the scenarios involving various potential user sub-intents are harder to handle than specific user intents, which need to be further investigated in the future.
 4.5 Efficiency Analysis of Ranking Algorithms
@@
 This phenomenon further proves the robustness of our ranker’s efficiency across diverse ranking settings.
-Due to limited space, we provide further analysis studies in Appendix C,  D and  E.
+Due to limited space, we provide further analysis studies in Appendix C,  D,  E,  F and  G.
 Figure 5: Efficiency experiments of different models.
@@
 Specifically, for the WikiPassageQA dataset, we prompt GPT-4 to generate appropriate and accurate sub-aspects and sub-answers for each data point to support our study, where sub-answers are split from the original long-form answers and sub-aspects are closely related to original questions and sub-answers.
-The content of the prompt is shown in Prompt E.
+The content of the prompt is shown in Prompt G.
+To ensure the quality of reformulated WikiPassageQA, the prompt of GPT-4 is decided via the following steps:
+(1) Human-Annotated Demonstrations: We first selected five data examples and manually annotated their sub-aspects and sub-answers.
+These human-annotated examples are used as demonstrations to prompt the GPT-4 to generate satisfied results.
+(2) Calibration with Human verification: Then, we randomly sampled 50 examples and used them as calibration examples.
+Specifically, we ceaselessly adjusted the content of the prompt based on the generation quality of these calibration examples until we thought the generation quality of these examples was satisfied.
+The satisfaction rate of human evaluation was 94%.
 The WikiAsp dataset is devised for generating aspect-based summaries of entities from 20 domains.
-Each piece of data is built from a Wikipedia article, consisting of various aspects of this article and aspect-based summaries, each target summary corresponds to an aspect.
+Each piece of data is built from a Wikipedia article, consisting of various aspects of this article and aspect-based summaries.
+Each target summary corresponds to an aspect.
 We follow Jiang et al.
 (2023) to convert it into open-domain QA settings and introduce some additional operations to make it suitable for our experimental settings.
-Firstly, samples with more than two aspects are retained and the remaining parts are removed to ensure the multifaceted nature of the experimental samples.
-Then, due to the expensive costs of experiments and the large amount of the whole dataset, we evenly sample subsets from each domain to build the experimental samples, and split training, validation, and test sets according to the ratio of 10:1:1.
+Firstly, samples with more than two aspects are retained, and the remaining parts are removed to ensure the multifaceted nature of the experimental samples.
+Then, due to the expensive costs of experiments and the large amount of the whole dataset, we evenly sample subsets from each domain to build the experimental samples and split training, validation, and test sets according to the ratio of 10:1:1.
 Finally, we insert the title of the original Wikipedia article into a template: “Generate a summary about title” to mimic the real question format, hence constructing the question of each sample data.
@@
 We concatenate these sub-answers to build the long-form answer for each sample.
-The statistical information of datasets is presented in Table 3.
+The statistical information of our datasets is presented in Table 2.
 Items WikiPassageQA
@@
 (2023) as our fixed retriever.
+All our baselines are optimized and evaluated by the same training, validation, and test datasets.
+Since the WikiAsp dataset has no relevance labels on documents, for each query,we view the documents whose matching score (the average of rouge scores evaluated by the ground truth answer) is higher than 0.5 as relevant documents to support the training of our baseline ranking models.
 Our experiments are conducted on the platform with four NVIDIA A100-SXM4-80GB GPUs.
 We will release our codes upon the acceptance of our study.
+Models WikiPassageQA WikiAsp
+MAP N1 N3 N5 N10 NCOM MAP N1 N3 N5 N10 NCOM
+Retriever .3201 .2476 .2933 .3498 .3951 .7006 .1780 .1003 .1678 .1953 .2294 .3926
+RankT5 (point) .5200 .4808 .5176 .5424 .5916 .7865 .3712 .3230 .3704 .3942 .4043 .4739
+RankT5 (pair) .3602 .2740 .3367 .3863 .4298 .7395 .3299 .2088 .3274 .3717 .3894 .5297
+LDIST (point) .5346 .5024 .5261 .5573 .5979 .7608 .4051 .3668 .4078 .4213 .4280 .5043
+LiT5 (list) .3756 .2813 .3661 .3968 .4412 .7465 .2757 .1696 .2742 .3142 .3318 .5442
+RAG-Fusion .2632 .1923 .2348 .2754 .3317 .5771 .2040 .1188 .1844 .2271 .2608 .4415
++RankT5 (point) .3962 .3510 .3831 .4238 .4641 .6543 .1591 .0681 .1411 .1742 .2177 .4610
++RankT5 (pair) .3946 .3029 .3669 .4202 .4758 .6717 .3297 .2376 .3249 .3583 .3869 .4992
++LDIST (point) .3746 .3245 .3654 .3951 .4385 .6370 .3345 .2987 .3241 .3505 .3733 .5138
++LiT5 (list) .3114 .2308 .2811 .3244 .3839 .6547 .2737 .1995 .2606 .3014 .3255 .5196
+BGM (list) .2614 .2476 .2480 .2626 .2864 .7988 .1382 .0727 .1152 .1497 .1921 .3887
+RichRAG (list) .5444 .5240 .5359 .5663 .6035 .8065 .4880 .4556 .4962 .5047 .5067 .8935
+Table 3: Overall ranking results. The best and second-best results are in bold and underlined, respectively.
+Appendix C Ranking Performance
+To prove the effectiveness of our proposed list-wise ranking module, we also evaluate the ranking modules’ performance of our method and baselines.
+We indicate the types of ranking algorithms, including point-wise and list-wise.
+We further implement a pairwise version of the RankT5 model by using the following training objective to optimize it:
+$\mathcal{L}=\!\!\!\!\!\!\!\!\!\!\!\!\!\!\sum_{d_{1},d_{2}\in\mathcal{D},R(d_{1})>R(d_{2})}\!\!\!\!\!\!\!\!\!\!\!\!\!\!\max(0,s(d_{2})-s(d_{1})+\gamma),\;\;\gamma=1,$
+where $\mathcal{D}$ denotes the candidate documents, $R(d)$ represents the relevance label of the document and $s(d)$ is the predicted score of the document.
+Specifically, we select some widely-used metrics, MAP and NDCG@k(k=1,3,5,10) to assess the ability to predict the document relevance of models.
+Furthermore, our model not only focuses on document relevance but also the comprehensiveness of the provided ranking lists.
+Thus, we propose a novel ranking metric, Normalized Comprehensiveness (NCOM), to assess the comprehensiveness of ranking lists, following the way to build the silver ranking targets (introduced in Section 3.4.1).
+This metric considers both the document relevance and the coverage of ranking lists on sub-aspects, hence reliably evaluating the comprehensiveness of ranking lists.
+Its calculation is presented below:
+Given a question $q$, its sub-answers $\{a_{1},...a_{n}\}$, a pool of candidate document $\mathcal{P}$, a ranking list generated by a ranking model, $L=[d^{1},...,d^{K}]$, and a silver ranking list, $L^{*}$, generated by Eq.
+8 in our paper, we calculate the generated and silver ranking lists’ comprehensiveness scores as follow:
+$COM(L)$ $=\sum_{t=1}^{K}\sum_{i=1}^{n}w^{t}_{i}\cdot\phi(d^{t},a_{i}),$ (13)
+$w^{t}_{i}$ $=1-\mathrm{Norm}_{i}(\max_{d\in L[:t]}\phi(d,a_{i}))$
+where $w^{t}_{i}$ denotes the importance of $i$-th sub-aspects at the $t$-th step (which is the same as the Equation 9 in our paper.), and $\phi(\cdot,\cdot)$ calculates the similarity between two sentences, i.e., average of Rouge-2, and Rouge-L scores.
+The final normalized comprehensiveness score is computed by normalizing $COM(L)$ by $COM(L^{*})$:
+$NCOM=COM(L)/COM(L^{*}).$
+The evaluation results are shown in Table 2.
+From the experimental results, we can find that our method outperforms all types of ranking algorithms on both datasets and all metrics, especially on NCOM.
+This phenomenon proves that our ranking model could consider the document’s relevance and list comprehensiveness simultaneously, hence providing comprehensive external knowledge and stimulating LLMs to generate better responses.
+We also notice that even though LiT5 and BGM are list-wise ranking models, their optimization does not explicitly consider the relationships among documents and, hence cannot provide satisfactory results.
+It also confirms the superiority of our proposed training algorithm.
 Figure 6: Trend of model performance as Top-K changes on both two dataset.
-Appendix C Analysis of LLMs’ Preferences
+Figure 7: Subset experiments with different sub-aspect amounts.
+Appendix D Impact of Sub-aspect Amount
+To test the generalization of our framework with different sub-aspects numbers, which represent different search scenarios, we further split the test dataset into different sub-sets according to the number of sub-aspects.
+Questions with a sub-aspect amount less than two are divided into the narrow set, questions with a sub-aspect amount less than four are divided into the middle set, and the remaining questions are divided into the broad set.
+The models’ performances on these subsets are shown in Figure 7.
+Evidently, our framework outperforms all baselines on all subsets.
+This result verifies the robustness of our framework with diverse search scenarios.
+Furthermore, we find that the overall results on the broad set are worse than the remaining two sets.
+This phenomenon implies that the scenarios involving various potential user sub-intents are harder to handle than specific user intents, which need to be further investigated in the future.
+Appendix E Analysis of LLMs’ Preferences
 We mentioned that in the system of RAG, the downstream users of IR systems are no longer humans, but LLMs.
@@
 It further confirms the importance of repetition with potential emphasis and denoising effect.
-Appendix D Impact of Number of External Documents
+Appendix F Impact of Number of External Documents
 To further investigate the impact of different numbers of external knowledge on RAG performance, we vary the value of K and conduct corresponding experiments on our model and a baseline model that directly treats the top-k retrieved documents as provided references for the generator.
@@
 The improvement trendlines also imply that with limited external references, our model could show better response performance due to better ranking abilities.
-Appendix E Case Study
-To validate that our proposed framework has the ability to provide rich and comprehensive responses for multi-faceted questions, we further demonstrate a case study to compare the generated responses of RichRAG and a RAG baseline that directly generates responses based on retrieved documents.
-In Box E, we provide the input question, its labeled sub-aspects, and predictions from different compared models.
+Figure 8: Comparison with and without repeated constraints
+Appendix G Case Study
+To validate that our proposed framework has the ability to provide rich and comprehensive responses for multi-faceted questions, we further demonstrate a case study to compare the generated responses of RichRAG and an RAG baseline that directly generates responses based on retrieved documents.
+In Box G, we provide the input question, its labeled sub-aspects, and predictions from different compared models.
 The parts of generated responses are highlighted by different colors to indicate the comprehensiveness of different results.
@@
 This result further proves the ability of RichRAG to identify and explicitly model the query’s sub-aspects, thereby providing satisfying ranking results to enhance the richness and comprehensiveness of generated responses.
-Figure 6: Comparison with and without repeated constraints
+This result also reveals that though LLMs have the potential ability to generate comprehensive answers, comprehensive ranking lists that correctly cover the query’s sub-aspects can better stimulate LLMs’ ability to generate comprehensive responses.
+This can also be verified by our quantitative experiments where our model exhibits the best performance among baselines.
 Box E: A case to show the richness and comprehensiveness of our model’s generated responses. Query: Generate a summary about Instant Mom.